Apart from being in the education sector, Richfield also has a fleet of cars available for rent. 

Statistics shows that the Mercedes AMG G63 is the most rented car among their fleet and one in four customers request this make. If 5 of today’s rentals are selected, what is the probability that 2 clients would have rented an AMG G63?


1
Expert's answer
2021-05-04T13:18:45-0400

"n = 5 \\

p = \frac{1}{4} = 0.25 \\

q = 1-p = 0.75 \\

P(X=2) = C^5_2(0.25)^2(0.75)^{5-2} \\

= \frac{5!}{2!(5-2)!} \times 0.25^2 \times 0.75^3 \\

= 10 \times 0.0625 \times 0.4218 \\

= 0.2636"